A closed surface which encloses a charge has an electric flux of 15.8 N.m^2/C. What is the magnitude of the charge?


1
Expert's answer
2022-03-08T08:57:03-0500

The electric flux through surface


"\Phi=\oint EdA=\frac{q}{\epsilon_0}"


Hence, the net charge

"q=\Phi \epsilon_0\\
=15.8*8.85*10^{-12}=1.4*10^{-10}\:\rm C"